OpenStudy (ilovemath14):

HELP ASAP!! HAVE TO FINISH THIS TODAY!! FAN & MEDAL!! DIVIDE 1. (b^2-13b+48) / (b-7) 2. (4p^2-29p-15) / (p-8) 3. (2r^3-7r^2-54r+35) / (r-7) 4. (m^3+3m^2+11m+2) / (m+1) 5. (a^2-7a+19) / (a-5) 6. (5r^2-15r+3) / (5r-5) 7. (80r^3-144r^2+164r-88) / (10r-8) 8. (b^3+b^2-43b+8) / (b-6)

OpenStudy (luckycoins888):

Factoring b2 - 13b + 48 The first term is, b2 its coefficient is 1 . The middle term is, -13b its coefficient is -13 . The last term, "the constant", is +48 Step-1 : Multiply the coefficient of the first term by the constant 1 • 48 = 48 Step-2 : Find two factors of 48 whose sum equals the coefficient of the middle term, which is -13 . -48 + -1 = -49 -24 + -2 = -26 -16 + -3 = -19 -12 + -4 = -16 -8 + -6 = -14 -6 + -8 = -14 For tidiness, printing of 14 lines which failed to find two such factors, was suppressed Observation : No two such factors can be found !! Conclusion : Trinomial can not be factored Polynomial Long Division : 1.2 Polynomial Long Division Dividing : b2-13b+48 ("Dividend") By : b-7 ("Divisor") dividend b2 - 13b + 48 - divisor * b1 b2 - 7b remainder - 6b + 48 - divisor * -6b0 - 6b + 42 remainder 6 Quotient : b-6 Remainder : 6 Final result : b2 - 13b + 48 ——————— b - 7 Note : you must simplify : b2 - 13b + 48 ——————— b - 7

OpenStudy (luckycoins888):

That is the first one.

OpenStudy (luckycoins888):

@ilovemath14 understand or no?

OpenStudy (ilovemath14):

Yes @luckycoins888

OpenStudy (luckycoins888):

kk

OpenStudy (luckycoins888):

medal pls xD

OpenStudy (ilovemath14):

than after that? I kinda know it @luckycoins888